# Multi-view Structural Convolution Network for Domain-Invariant Point Cloud Recognition of Autonomous Vehicles

### v3

@@ Abstract
Point cloud representation has recently become a research hotspot in the field of computer vision and has been utilized for autonomous vehicles. However, adapting deep learning networks for point cloud data recognition is challenging due to the variability in datasets and sensor technologies. This variability underscores the necessity for adaptive techniques to maintain accuracy under different conditions. In this paper, we present the Multi-View Structural Convolution Network (MSCN) designed for domain-invariant point cloud recognition. MSCN comprises Structural Convolution Layers (SCL) that extract local context geometric features from point clouds and Structural Aggregation Layers (SAL) that extract and aggregate both local and overall context features from point clouds. Additionally, our MSCN enhances feature representation robustness by training with unseen domain point clouds derived from source domain point clouds. This method acquires domain-invariant features and exhibits robust, consistent performance across various point cloud datasets, ensuring compatibility with diverse sensor configurations without the need for parameter adjustments. This highlights MSCN’s potential to significantly improve the reliability and domain invariant features in different environments. Our code is available at https://github.com/MLMLab/MSCN.
keywords:
@@
Point cloud-based Autonomous System, LiDAR Point Cloud Classification, Domain-Invariant Feature Representation, Unseen Domain Generation, Domain Generalization.
journal: Expert Systems with Applications
[…]
[a]organization=Department of Civil Engineering, University of Central Florida,postcode=32816, state=Florida, country=USA
[…]
[b]organization=Department of Data, Networks and AI, Ajou University,postcode=16499, state=Suwon, country=South Korea
[…]
[c]organization=Department of Mobility Engineering and Department of Artificial Intelligence, Ajou University,postcode=16499, state=Suwon, country=South Korea
1 Introduction
Point cloud sensors and data have gained significant attention and are utilized in various fields due to their potential to enhance perception capabilities in advanced systems and robotics through accurate point distance measurements. The point cloud data is generated by Light Detection and Ranging (LiDAR) sensors, which produce high-resolution 3D spatial data and offer accurate measurements of the surrounding environment. These point clouds are crucial for various applications, including autonomous vehicles, robotics, and smart infrastructure systems, where accurate environmental understanding is essential.
Despite considerable advances in 2D image-based object recognition, the transition to 3D object recognition from LiDAR-based point clouds remains a largely underexplored area. This is primarily due to the unstructured and unordered nature of point clouds, which introduces significant challenges for analysis and interpretation. Unlike 2D images, point clouds lack a clear grid structure, making traditional methods less effective. Overcoming these obstacles is critical for improving the accuracy and reliability of 3D perception systems, which rely heavily on efficient classification techniques.
Figure 1: Illustration of self-driving vehicle recognition scenarios from cross-source point clouds. The properties of the point cloud vary according to the channels in the LiDAR sensor; fewer channels result in sparser point clouds, while more channels create denser point clouds. Additionally, the characteristics of the point cloud differ not only based on the region where the training data was acquired but also between those obtained from simulations and those collected in the real world, which have very different characteristics. Therefore, to maintain consistent recognition capabilities of self-driving cars, robust learning across domains, including changes in LiDAR sensors and areas (real to real and sim to real), is essential.
Moreover, ensuring the robust and consistent trained model is essential for perception, localization and mapping in autonomous vehicles. As depicted in Figure 1, the quality of point clouds obtained through LiDAR sensors heavily depends on the sensor’s resolution. However, LiDAR sensors vary significantly in resolution depending on the manufacturer and its model, leading to substantial differences in the quality of point clouds used for recognition models. Furthermore, real-world objects exhibit significant variation based on its installed location or multi-sensor usage, adding another layer of complexity. These disparities in real-world point clouds have different point cloud patterns and resolutions and result in misjudgments on perception in self-driving vehicles. For example, during the development phase, self-driving vehicles may be trained using high-resolution LiDAR sensors to produce high-quality point clouds. However, in practical deployment scenarios, commercial vehicles may be equipped with low-resolution LiDAR sensors due to cost or other constraints. This disparity can lead to severe performance gaps between the development and deployment environments. Furthermore, when self-driving vehicles are deployed in new unvisited regions or environments during training, models can experience a significant drop in performance[30].
Despite these challenges, recent studies have largely overlooked the domain gap between training and deployment scenarios. While these models may perform well on point clouds with geometric properties similar to the training dataset, their performance often deteriorates in diverse and unseen real-world scenarios. Thus, existing research has failed to address the variability and inconsistencies inherent in real-world point clouds, leaving a critical domain gap in deploying models for various domains in different self-driving platforms.
To process point cloud data from LiDAR, approaches such as point-based [1], [2], [4] and graph-based [3] have been proposed. PointNet [1] processes a point cloud using a neural network to directly consume raw point cloud data, extracting global features through symmetric functions and max pooling to achieve permutation invariance. However, PointNet does not take into account the local structural feature of the point cloud, which may lead to overfit to certain dataset. PointNet++ [2] learns the hierarchical features of point cloud data by stacking multiple learning stages, allowing it to capture detailed local geometric information, but the high computational complexity of this network limits its application to real-world scenarios. PointMLP [4] allows the structural characteristics of point clouds to be extracted using only a simple MLP with residual connections and a geometric affine module without a local geometric extractor, achieving state-of-the-art results towards the ModelNet40 and ScanObjectNN datasets. PointMLP can represent structural features in the overall context of a 3D object, but because it overlooks structural features in the local context, there may be significant performance degradation for outdoor datasets with occlusion or data distribution variance. Dynamic Graph CNN (DGCNN) [3] enhances the understanding of 3D geometric information by using graph-based convolution networks and a k-nearest neighbor strategy to form dynamic graphs, improving local feature comprehension. However, the complexity of DGCNN introduces computational challenges and relies heavily on the precision of input graphs, necessitating innovative solutions for robust feature extraction.
To describe the local information of 3D data more in-depth, some research works [5], [30] utilize deformable 3D kernels that are learnable during the training stage. 3D Graph Convolutional Networks (3D-GCN) [5], distinguished by their incorporation of geometric values and adjustment of network parameters through directional analysis, employ deformable graph kernels to adeptly manage translation, scale, and z-axis rotation. However, this approach cannot properly explain the distances between points in a point cloud, potentially leading to performance degradation in case of domain change. This is mainly because point clouds acquired in the real world have very different spacings between points depending on the dataset. 3D Structural Convolution Network (3D-SCN) [30] identifies important geometric structures of a point cloud by analyzing the relationship of points between neighborhoods from the standpoint of both distance and direction, enabling domain-invariant point cloud recognition. Nevertheless, because 3D-SCN only trains the local geometric information of a single point cloud, it overlooks the structural information in the overall context and may miss important structural properties of the object.
@@ 1 Introduction
In this paper, we propose a novel deep learning model named Multi-view Structural Convolution Network (MSCN), which can extract features of 3D point clouds that are robust to domain changes in order to respond to arbitrary domain changes that frequently occur in point cloud recognition scenarios. The MSCN consists of a Structural Convolution Layer (SCL) and a Structural Aggregation Layer (SAL). In the SCL, structural information between points within the point cloud is captured in a local context, while in the SAL, structural information of the point cloud is captured in a global context and embedded into each point across multiple scales. By analyzing the structure of the point cloud from multiple perspectives, our MSCN can achieve highly sophisticated structural feature representation, ensuring stable performance despite domain changes.
@@
In addition, point clouds are inherently irregular and exhibit significant variability depending on the LiDAR resolution and geographic location. To address these real-world irregularities and improve the robustness of the model, we propose a novel approach for generating unseen point clouds based on source point clouds. Inspired by the Progressive Domain Expansion Network [14] proposed for single-domain generalization in 2D images, we extend this concept to 3D point cloud scenarios. By repeatedly generating virtual point clouds of new domains and training the model with these alongside the existing source point clouds, the model learns to generalize effectively to unseen point clouds encountered in real-world scenarios. This process enables the model to achieve greater invariance to domain changes and maintain consistent performance in various real-world conditions.
The main contributions of this study can be summarized in three folds as follows:
  • 1.
@@
    We propose a point cloud domain explansion framework by modifying the 2D Progressive Domain Expansion Network (PDEN)[14] for effectively generating unseen domain point clouds, utilizing these generated virtual point clouds to train the model, thereby enhancing the MSCN’s robustness to domain changes.
  • 3.
    We propose a synthetic point cloud dataset to evaluate the consistent performance of our model when transitioning from simulation to the real world. Furthermore, we demonstrate the data-invariant performance of the proposed MSCN (Multi-view Structural Convolution Network) on this synthetic dataset as well as on various LiDAR-based real-world point cloud datasets, including the nuScenes dataset (32 channels, United States), the KITTI dataset (64 channels, Germany), and the PanKyo dataset (128 channels, South Korea). These demonstrations highlight the robust adaptability of our model to different sensor configurations and locations.
@@
The subsequent sections of this paper are organized as follows. Section 2 presents a comprehensive literature review of 3D point cloud encoding algorithms, domain adaptation methods, and domain generalization methods. Section 3 illuminates the proposed MSCN architecture. Section 4 shows a framework for progressively generating point clouds of unseen domains from a single domain point cloud and learning a feature extractor with the generated point clouds. In Section 5, the paper presents an evaluation of the proposed network using various point cloud datasets acquired from synthetic datasets, real-world indoor dataset, and real-world self-driving vehicle datasets. Finally, Section 6 concludes the paper.
2 related works
2.1 3D point cloud deep learning networks: Encoding perspectives
2.1.1 Multi-view-based methods
@@
Considering that a collection of 2D views can offer significant information for 3D shape recognition compared to traditional manual feature extraction, several studies [15], [16], [17], [18] have explored recognizing 3D shapes from collections of rendered 2D images. Inspired by these works, Hamdi et al. [19] proposed the Multi-View Transition Network (MVTN). This network includes a differentiable module that predicts the optimal viewpoint for a task-specific multi-view network, aiming to enhance 3D shape recognition by addressing limitations of previous multi-view methods. However, multi-view approaches still struggle to fully exploit the geometric relationships inherent in 3D data and extend their applicability to other 3D tasks, such as segmentation and reconstruction.
2.1.2 Voxel-based methods
@@
To manage the unstructured nature of point clouds, some studies [20], [21], [22] have used voxel grids or voxelized 3D shapes as model inputs, enabling the application of standard 3D CNNs. Researchers like Riegler et al. [23] and Wang et al. [24] have employed octree structures instead of voxels to improve computational efficiency and reduce memory consumption. Despite their focus on the relationships within 3D data and the grouping of point clouds, voxel-based methods often suffer from low classification efficiency due to the inherent sparsity and incomplete information in point clouds.
2.1.3 Point cloud-based methods
@@
2.2 3D point cloud deep learning networks: Local aggregation
2.2.1 Point-by-point methods
@@
2.2.2 Graph-based methods
@@
2.3 Domain adaptation and Domain generalization algorithms
2.3.1 Domain adaptation
@@
Recent studies [6], [7] have demonstrated that deep learning models are prone to significant performance degradation when evaluated on out-of-distribution (OOD) datasets, even with minor variations in the data generation process. These studies underscore the limited generalization capacity of deep learning models, which can severely impact their performance in real-world applications. Domain adaptation (DA) [8], [9], [10] has emerged as a strategy to tackle the OOD data challenge by utilizing a subset of data from the target domain. This approach enables a model trained on a source domain with different distributions to adapt to the target domain. The fundamental concept of DA is that, despite differences between source and target domains, common underlying structures or patterns can be leveraged to enhance model performance in the target domain. However, DA assumes the availability of either labeled or unlabeled data from the target domain for model adaptation, a condition not always met in real-world scenarios. Acquiring labeled data from the target domain can be prohibitively time-consuming, expensive, or impractical due to ethical or legal constraints. Additionally, the target domain’s data distribution may remain unclear, complicating the application of conventional DA methods that rely on understanding this distribution.
2.3.2 Single domain generalization
@@
To address the challenges associated with domain adaptation (DA), many domain generalization (DG) studies [11], [12], [13], [14] particularly those focusing on single domain generalization, have been proposed. Single domain generalization is a particularly challenging form of model generalization where models are trained on a single domain and tested on unseen domains. A promising approach involves learning cross-domain invariant representations by expanding the coverage of the training domain. These methods typically synthesize additional samples in feature space to broaden the data distribution in the training set. Zhang et al. [11] observed that differences in medical images mainly arise in three aspects: image quality, image appearance, and spatial configuration. They augment these aspects through data augmentation. However, this method requires selecting the appropriate type and magnitude of augmentation based on the target domain, which can be challenging for other tasks. Volpi et al. [12] and Qiao [13] synthesize additional data through adversarial learning to enhance model robustness. Nevertheless, this approach has limitations: the types of augmentation are relatively simple, and using too many adversarial examples during training can degrade classifier performance. To mitigate these issues, Li et al. [14] proposed progressively generating multiple domains and learning a feature extractor using both source and generated data. The model employs contrastive learning to develop domain-invariant representations, effectively clustering each class to facilitate better decision boundaries and improve generalization. While this approach equips models to perform well in real-world scenarios, particularly on previously unseen domains, it is currently only applicable to domain generalization tasks involving well-structured 2D images.
Figure 2: The Multi-view Structural Convolution Network (MSCN) architecture for 3D point cloud classification comprises several components. (a) The MSCN includes feature extraction layers, specifically Structural Convolution Layers (SCL) and Structural Aggregation Layers (SAL), as well as Global Max Pooling and a multi-layer perceptron (MLP) for classification. (b) SCL is engineered to extract local features from each point. (c) The SAL is designed to combine local context features with overall context features.
3 Multi-view Strucutral Convolution Network(MSCN)
@@
3.1 Structural Convolution Layer(SCL)
@@
$Conv_{dir}(R^{M}_{n},K^{S}_{dir})=\langle F(p_{n}),w(b^{c}_{dir})\rangle+\sum\limits_{s=1}^{S}\max\limits_{m\in(1,M)}sim(p_{m},b^{s}_{dir})$ (4)
@@
$Conv_{dist}(R^{M}_{n},K^{S}_{dist})=\sum_{s=1}^{S}w_{dist}\times\max\limits_{m\in(1,M)}(||d_{m,n}||)$ (6)
@@
$SCL(D_{i-1},D_{i})=MLP(Conv_{dir}(R^{M}_{n},K^{S}_{dir})\oplus Conv_{dist}(R^{M}_{n},K^{S}_{dist}))$ (7)
3.2 Structural Aggregation Layer (SAL)
SAL $(D_{i-1},D_{i})$ is specially designed to capture the structural information of a point cloud from both local and global perspectives. As shown in Figure 2 (c), SAL$(D_{i-1},D_{i})$ receives a point cloud from the previous layer, where all points have high-dimensional local features $D_{i-1}$ as input. This layer first performs neighborhood sampling to construct the receptive fields and then applies channel-wise max-pooling within these receptive fields to aggregate features $F(p)$, $\forall p\in R^{M}_{n}$, resulting in a coarse to fine feature representation at the local level. Next, a subset of $P$ is downsampled using a predetermined sampling rate $r$. These local view aggregation processes can be formulated as:

### v4

# MSCN: Multi-view Structural Convolution Network for Domain-Invariant Point Cloud Recognition of Autonomous Vehicles

@@ Abstract
Although LiDAR sensors have become indispensable for autonomous vehicles (AVs) due to their ability to provide accurate 3D scene understanding and robust perception under adverse weather conditions, the properties of LiDAR point clouds vary widely across sensor configurations and data acquisition domains, leading to severe performance degradation when models are transferred between heterogeneous sensors or from simulation to the real world. To address this challenge, we propose the Multi-view Structural Convolution Network (MSCN), a novel architecture designed to achieve domain-invariant recognition across diverse LiDAR configurations and environments. MSCN comprises Structural Convolution Layers (SCL) that extract local context geometric features from point clouds and Structural Aggregation Layers (SAL) that extract and aggregate both local and overall context features from point clouds. Furthermore, we incorporate an unseen domain generation strategy to mitigate domain gaps during training. Extensive experiments demonstrate that MSCN consistently outperforms state-of-the-art point cloud classification methods across all domain change scenarios. These results highlight MSCN as a scalable solution for deploying LiDAR-based perception systems of AVs. Our code is available at https://github.com/MLMLab/MSCN.
Index Terms:
@@
Point cloud-based Autonomous Vehicle System, LiDAR Point Cloud Classification, Domain-Invariant Feature Representation, Unseen Domain Generation, Domain Generalization.
I Introduction
Autonomous vehicles (AVs) are envisioned as a transformative solution to major transportation challenges, including the mitigation of traffic accidents [1, 2], congestion alleviation [3, 4], and improved efficiency in transportation systems [5]. Reliable perception is central to enabling these benefits, as AVs must accurately recognize and interpret their surroundings to ensure safety and efficiency. In this context, Light Detection and Ranging (LiDAR) has emerged as a particularly valuable sensor because it provides rich geometrical information, such as depth and structural details, that complements traditional camera-based perception. LiDAR-based point clouds provide accurate geometric representations of traffic environments, offering precise depth and structural information that is indispensable for robust AV perception and intelligent transportation systems (ITS).
With the advantages of LiDAR sensors, recent studies have increasingly leveraged LiDAR-centered perception within AVs for accurate 3D scene understanding with rich geometrical information and for robust perception under adverse weather conditions where camera-based systems often fail [8, 9, 10, 11]. For example, recent work demonstrates that LiDAR maintains reliable detection and localization performance under adverse weather conditions such as rain, fog, and snow, highlighting its robustness where camera-based systems fail [8]. Complementing this, Bhardwaj et al. [9] propose a memristive associative learning circuit for fault-tolerant multi-sensor fusion, showing that incorporating LiDAR with other modalities ensures robust perception even when certain sensors degrade or fail in harsh environments.
Figure 1: Illustration of LiDAR-based recognition’s three key challenges, including different sensor configurations, geographic locations, and sim-to-real gap.
Despite the advances of LiDAR-based studies, prior research has largely overlooked the importance of achieving robust perception across heterogeneous LiDAR sensors. In practice, as illustrated in Figure 1, the properties of point clouds vary significantly depending on both the sensor configuration and the data acquisition domain. First, different sensor configurations, such as variations in the number of LiDAR channels, cause substantial discrepancies in point cloud density and resolution. When models are trained on high-resolution LiDARs but deployed on cost-efficient low-resolution LiDARs, they often experience severe performance degradation. Since LiDAR configurations vary widely across manufacturers and even across research platforms, this mismatch presents a critical barrier to the scalable adoption of autonomous driving systems. Second, point cloud characteristics vary across geographic locations, as object appearances and structural patterns differ in distinct environments. Without accounting for such location-specific variations, autonomous vehicles risk misinterpreting unfamiliar object shapes, which can lead to failures in recognition and ultimately compromise driving safety. Moreover, the widely recognized sim-to-real gap poses another major challenge: while simulation environments enable the generation of abundant data and diverse accident scenarios, the point clouds they produce are typically uniform and noise-free. By contrast, real-world LiDAR data are often sparse and irregular, resulting in a pronounced distributional discrepancy that limits the generalization of models trained solely on simulation data.
@@ I Introduction
To address these challenges, we propose a novel deep learning model named Multi-view Structural Convolution Network (MSCN), which can extract features of 3D point clouds that are robust to domain changes in order to respond to arbitrary domain changes that frequently occur in point cloud recognition scenarios. The MSCN consists of a Structural Convolution Layer (SCL) and a Structural Aggregation Layer (SAL). In the SCL, structural information between points within the point cloud is captured in a local context, while in the SAL, structural information of the point cloud is captured in a global context and embedded into each point across multiple scales. By analyzing the structure of the point cloud from multiple perspectives, our MSCN can achieve highly sophisticated structural feature representation, ensuring stable performance despite domain changes.
@@
In addition, point clouds are inherently irregular and exhibit significant variability depending on the LiDAR resolution and geographic location. To address these real-world irregularities and further improve the robustness of the model, we propose a novel approach for generating unseen point clouds based on source point clouds. Inspired by the Progressive Domain Expansion Network proposed for single-domain generalization in 2D images [25], we extend this concept to 3D point cloud scenarios. By repeatedly generating virtual point clouds of new domains and training the model with these alongside the existing source point clouds, the model learns to generalize effectively to unseen point clouds encountered in real-world scenarios. This process enables the model to achieve greater invariance to domain changes and maintain consistent performance in various real-world conditions.
The main contributions of this study can be summarized as follows:
@@
  • We propose a point cloud domain expansion framework by modifying the 2D Progressive Domain Expansion Network (PDEN)[25] for effectively generating unseen domain point clouds, utilizing these generated virtual point clouds to train the model, thereby enhancing the MSCN’s robustness to domain changes.
  • We introduce a synthetic point cloud dataset to evaluate the consistent performance of our model in sim-to-real transitions. This dataset provides a controlled environment to validate the robustness and effectiveness of MSCN under simulated conditions.
  • We design progressively harder experimental settings to rigorously evaluate the domain-invariant recognition properties of MSCN: (1) uniform synthetic point clouds with controlled geometric perturbations, (2) indoor real-world point clouds with geometric variations, (3) autonomous vehicle datasets with varying LiDAR channel configurations and geographic locations, (4) sim-to-real domain shifts from synthetic to real driving datasets, and (5) combined challenges of channel change, geographic variation, and geometric perturbations in autonomous vehicle datasets. These progressively harder settings demonstrate MSCN’s robust geometric-invariant recognition properties, which are essential for ensuring reliable perception in AVs under diverse deployment conditions.
@@
The subsequent sections of this paper are organized as follows. Section II presents a comprehensive literature review of 3D point cloud encoding algorithms, domain adaptation methods, and domain generalization methods. Section III illuminates the proposed MSCN architecture. Section IV shows a framework for progressively generating point clouds of unseen domains from a single domain point cloud and learning a feature extractor with the generated point clouds. In Section  V, the paper presents an evaluation of the proposed network using various point cloud datasets acquired from synthetic datasets, real-world indoor dataset, and real-world self-driving vehicle datasets. Finally, Section VI concludes the paper.
II Literature Review
II-A 3D point cloud deep learning networks: Encoding perspectives
II-A1 Multi-view-based methods
@@
Considering that a collection of 2D views can offer significant information for 3D shape recognition compared to traditional manual feature extraction, several studies have explored recognizing 3D shapes from collections of rendered 2D images [26], [27], [28], [29]. Inspired by these works, Hamdi et al. proposed the Multi-View Transition Network (MVTN) [30]. This network includes a differentiable module that predicts the optimal viewpoint for a task-specific multi-view network, aiming to enhance 3D shape recognition by addressing limitations of previous multi-view methods. However, multi-view approaches still struggle to fully exploit the geometric relationships inherent in 3D data and extend their applicability to other 3D tasks, such as segmentation and reconstruction.
II-A2 Voxel-based methods
@@
To manage the unstructured nature of point clouds, some studies have used voxel grids or voxelized 3D shapes as model inputs, enabling the application of standard 3D CNNs [31], [32], [33]. Researchers like Riegler et al. [34] and Wang et al. [35] have employed octree structures instead of voxels to improve computational efficiency and reduce memory consumption. Despite their focus on the relationships within 3D data and the grouping of point clouds, voxel-based methods often suffer from low classification efficiency due to the inherent sparsity and incomplete information in point clouds.
II-A3 Point cloud-based methods
@@
II-B 3D point cloud deep learning networks: Local aggregation
II-B1 Point-by-point methods
@@
II-B2 Graph-based methods
@@
II-C Domain adaptation and Domain generalization algorithms
II-C1 Domain adaptation
@@
Recent studies have demonstrated that deep learning models are prone to significant performance degradation when evaluated on out-of-distribution (OOD) datasets, even with minor variations in the data generation process [17], [18]. These studies underscore the limited generalization capacity of deep learning models, which can severely impact their performance in real-world applications. Domain adaptation (DA) has emerged as a strategy to tackle the OOD data challenge by utilizing a subset of data from the target domain [19], [20], [21]. This approach enables a model trained on a source domain with different distributions to adapt to the target domain. The fundamental concept of DA is that, despite differences between source and target domains, common underlying structures or patterns can be leveraged to enhance model performance in the target domain. However, DA assumes the availability of either labeled or unlabeled data from the target domain for model adaptation, a condition not always met in real-world scenarios. Acquiring labeled data from the target domain can be prohibitively time-consuming, expensive, or impractical due to ethical or legal constraints. Additionally, the target domain’s data distribution may remain unclear, complicating the application of conventional DA methods that rely on understanding this distribution.
II-C2 Single domain generalization
@@
To address the challenges associated with DA, many domain generalization (DG) studies, particularly those focusing on single domain generalization, have been proposed [22], [23], [24], [25]. Single domain generalization is a particularly challenging form of model generalization where models are trained on a single domain and tested on unseen domains. A promising approach involves learning cross-domain invariant representations by expanding the coverage of the training domain. These methods typically synthesize additional samples in feature space to broaden the data distribution in the training set. Zhang et al. [22] observed that differences in medical images mainly arise in three aspects: image quality, image appearance, and spatial configuration. They augment these aspects through data augmentation. However, this method requires selecting the appropriate type and magnitude of augmentation based on the target domain, which can be challenging for other tasks. Volpi et al. [23] and Qiao et al. [24] synthesize additional data through adversarial learning to enhance model robustness. Nevertheless, this approach has limitations: the types of augmentation are relatively simple, and using too many adversarial examples during training can degrade classifier performance. To mitigate these issues, Li et al. [25] proposed progressively generating multiple domains and learning a feature extractor using both source and generated data. The model employs contrastive learning to develop domain-invariant representations, effectively clustering each class to facilitate better decision boundaries and improve generalization. While this approach equips models to perform well in real-world scenarios, particularly on previously unseen domains, it is currently only applicable to domain generalization tasks involving well-structured 2D images.
III Multi-view Structural Convolution Network (MSCN)
@@
III-A Structural Convolution Layer (SCL)
@@
Figure 2: Illustration of the Multi-view Structural Convolution Network (MSCN) architecture, including (a) the classification architecture, (b) the Structural Convolution Layer, and (c) the Structural Aggregation Layer.
@@
$\begin{split}\mathrm{Conv}_{\mathrm{dir}}(R^{M}_{n},K^{S}_{\mathrm{dir}})&=\left\langle F(p_{n}),\,w\!\left(b^{c}_{\mathrm{dir}}\right)\right\rangle\\ &\quad+\sum_{s=1}^{S}\max_{m\in(1,M)}sim\!\left(p_{m},\,b^{s}_{dir}\right)\end{split}$ (4)
where the symbol $\langle\cdot\rangle$ denotes the inner product operation, and the function simsim calculates the inner product between the features $F(p_{m})$ and the directional weights $w(b^{s}_{dir})$, utilizing cosine similarity [16] to define this interaction:
@@
[…]
@@
$SCL(D_{i-1},D_{i})$
$\hskip-68.00012pt=\mathrm{MLP}\!\Big{(}\operatorname{Conv}_{\mathrm{dir}}(R^{M}_{n},K^{S}_{\mathrm{dir}})\oplus\operatorname{Conv}_{\mathrm{dist}}(R^{M}_{n},K^{S}_{\mathrm{dist}})\Big{)}$ (7)
III-B Structural Aggregation Layer (SAL)
[…]